# Self-Controlled Dynamic Expansion Model for Continual Learning

### v1

@@ 4.3. Ablation study
Analysis of W-D Constraint (FDC). To evaluate the effect of feature alignment via Wasserstein Distance (W-D), we randomly selected 20 classes from TinyImageNet and visualized the feature representations using t-SNE and UMAP. fig. 3.(a) and fig. 3.(b) clearly demonstrate that features refined by the W-D constraint remain consistently closer to the Baseline distribution—i.e., the output of the frozen, pretrained backbone—compared to the unconstrained finetuned version. This observation suggests that the W-D constraint successfully preserves the semantic geometry of the pretrained feature space while allowing adaptation to new tasks.
The histograms in (c) and (d) further confirm this effect quantitatively: SCDEM achieves lower cosine distances to the baseline than the finetuned model (0.23 vs. 0.36 and 0.19 vs. 0.46, respectively), indicating a significantly smaller deviation from the original representation. From a modeling perspective, the resulting fused features are regularized against their historical counterparts using a differentiable Sinkhorn-based approximation of Wasserstein-2 distance. This formulation encourages the evolving features to align in both global and local geometry, providing a smooth inductive bias that mitigates representational drift without requiring access to previous data. Hence, the W-D constraint acts as a soft structural prior, preserving generality and stability in our task setting. We provide the additional ablation results in Appendix-C from SM.

### v2

@@ 4.3. Ablation study
Analysis of W-D Constraint (FDC). To evaluate the impact of feature alignment using Wasserstein Distance (W-D), we randomly selected 20 classes from TinyImageNet and visualized feature distributions via t-SNE and UMAP. As illustrated in fig. 3.(a) and (b), features constrained by W-D remain significantly closer to the Baseline distribution—i.e., the output of the frozen pretrained backbone—compared to the unconstrained fine-tuned version. This suggests that W-D helps preserve the semantic geometry of the original feature space while enabling adaptation to new tasks.
The histograms in (c) and (d) further quantify this effect: SCDEM achieves notably lower cosine distances to the Baseline (0.23 vs. 0.36 and 0.19 vs. 0.46), reflecting a smaller deviation from the pretrained representations. From a modeling perspective, the fused features are softly regularized toward their historical counterparts, encouraging geometric alignment in both global and local structure. This alignment acts as a structural prior that supports stable yet flexible representation learning across tasks. Additional results are provided in Appendix-C from SM.